OpenStudy (anonymous):

Liza kept the following records of her utility bills for 12 months: $29, $24, $32, $28, $45, $30, $29, $62, $40, $39, $33, $29 (a) Find the mean of Liza's monthly utility bill. (b) Find the median of Liza's monthly utility bill. (c) Is the mean or median a more useful representative of Liza's monthly utility bills?

OpenStudy (anonymous):

mean= sum of all the numbers/12 for median arrange the values in ascending order, median is the sum of two middle values divided by 2

OpenStudy (anonymous):

So the mean is 35?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

24, 28, 29, 29, 29, 30, 32, 33, 29, 40, 45, 62 median = 30 + 32/2=62/2=31

OpenStudy (anonymous):

and the median is 24,28,29,29,29,30,32,33,39,40,45?=31?

OpenStudy (anonymous):

Is the mean or median a more useful representative of Liza's monthly utility bills?

OpenStudy (anonymous):

i guess median is better representation

OpenStudy (anonymous):

(a) Find the mean of Liza's monthly utility bill. The mean is the sum of the bills divided by the number of the bills \[\frac{29+24+32+28+45+30+29+62+40+39+33+29}{12}=35 \] (b) Find the median of Liza's monthly utility bill. The Median is the 'middle value' in the list. When the totals of the list are odd, the median is the middle entry in the list after sorting the list into increasing order. When the totals of the list are even, the median is equal to the sum of the two middle (after sorting the list into increasing order) numbers divided by two. Sort[{29, 24, 32, 28, 45, 30, 29, 62, 40, 39, 33, 29}] = {24, 28, 29, 29, 29, 30, 32, 33, 39, 40, 45, 62} There are 12 bills. 12 is an even number. The middle two numbers are 30 and 32. In effect, and in this case, the median is the mean of 30 and 32 which is 31. (c) Is the mean or median a more useful representative of Liza's monthly utility bills? I personally have no opinion. I don't recall ever computing the median in everyday life. Flip a coin or do some research or determine what answer will make the teacher happy. The definition of median came from the following web site: http://math.about.com/od/statistics/a/MeanMedian.htm